OpenStudy (anonymous):

The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (anonymous):

length*width = area.... dividing both sides by length we get W=A/L width =(x2-2x-15)/x+3 factorizing the numerator we get (x+3)(x-5)/(x+3) =x-5 so width is x-5

OpenStudy (anonymous):

thank you!